40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14400     Entry time: Tue Jan 15 15:27:36 2019
Author: gautam 
Type: Update 
Category: General 
Subject: Lasers and other stuff turned back on 

VEA is now a laser hazard area as usual, several 1064nm lasers in the lab have been turned back on. Apart from this

  • the IFR was reset to the nominal modulation settings of +13dBm output at 11.066209 MHz (this has to be done manually following each power failure).
  • The temeprature control unit for the EY doubling oven PID control was turned back on.
  • The EY Oplev HeNe was turned back on.
  • EY green PZT HV Kepco was turned back on.
ELOG V3.1.3-